# Quantum simulating continuum field theories with large-spin lattice models

### v1

@@ E.2 $\beta^{2}$-renormalization and $\kappa$-dependence
To engineer the spatial derivative term $[\partial_{x}\hat{\varphi}(x)]^{2}$ in the sG model, we need to Taylor expand the cosine term $-V_{\text{nn}}^{\prime}/2\cdot\cos(\varphi_{i}-\varphi_{i+1})=-V_{\text{nn}}^{\prime}/2\cdot\cos(\delta\varphi_{i})$ of the lattice QFT Hamiltonian Eq. (38), which arises from the dipole-dipole interaction term $-\frac{1}{4}\left(\hat{J}_{+}^{(i)}\hat{J}_{-}^{(j)}+\mathrm{H.c.}\right)$. We thus obtain
@@
where we need to neglect higher-order term $(n\geq 2)$ to retrieve the sG model. Upon multiplying the Hamiltonian with $\beta^{2}/(2\chi\kappa^{2}a)$, mapping $\kappa\hat{\varphi}_{\text{latt}}\rightarrow\beta\hat{\varphi}_{\text{cont}}$, and using the relation $V_{\text{nn}}^{\prime}/\chi=4\kappa^{4}/\beta^{4}$, as we do in App. B to obtain the continuum sG model, the neglected terms become
@@ E.3 $M_{0}^{\prime}$ renormalization
E.4 Lukyanov-Zamolodchikov conjecture
@@
with $K=\beta^{2}/8\pi$ and is valid for $\beta^{2}<8\pi$, $|\Re\gamma|<1/(2\beta)$. For $\gamma=\sqrt{K}=\beta/\sqrt{8\pi}$ the above expression agrees with the theoretical prediction for the vertex operator of the sG model Eq. (14) [31].
Appendix F Semi-classical soliton preparation and dynamics
In this Appendix we provide additional details regarding the preparation of the semi-classical soliton, as well as its dynamics. This Appendix is structured as follows: in App. F.1 we discuss the prepared soliton state, and we show that the two-point correlators remain untouched, in App. F.2 resp. App. F.3 we focus on the time evolution of the static resp. moving soliton, in App. F.4 we discuss in more detail the semi-classical model, while in App. F.5 we illustrate the relation between the $zz$-correlator of the quantum ground state and the standard deviation of the velocity of the semi-classical model.
F.1 State preparation and correlation functions
@@
In Fig. 12(a) we plot the absolute value of the vertex operator $|\langle\hat{J}_{+}^{(i)}\rangle/\sqrt{J(J+1)}|$, which displays a perfect overlap between the ground state of the quantum sG model and the prepared static soliton, while for the moving soliton a small dip emerges in the center of the chain, right in the middle of the transition region. We attribute these deviations to the non-commutativity of the two sets of unitary rotations. Moreover, note that the absolute value of the vertex operator $|\langle e^{i\beta\hat{\varphi}}\rangle|$, i.e., the normalized “classical” spin length, is never 1 [see Fig. 4(c)]. Therefore, when determining the rotation angle $\theta^{(i)}$ of the gate $\hat{U}_{\text{m}}^{(i)}$ for the momentum imprinting, we need to take this local value into account, such that the magnitude of the momentum expectation value, which is proportional to the projection of the spin on the $z$-axis $|\langle\hat{\pi}\rangle|\propto|\langle\hat{J}_{z}\rangle|$, agrees with the theoretical prediction [see Fig. 12(b)].
@@
Figure 12: One- and two-point correlation functions of the prepared soliton state. Comparison of selected expectation values for different initial states: the solid black line represents the ground state of the quantum sG model, the blue circles the semi-classical static soliton $v=0$, the orange dots the semi-classical moving soliton at $v=0.5$, the dashed black line the theoretical prediction for the momentum of the classical moving soliton. (a) Absolute value of the expectation value of the vertex operator $|\langle\hat{J}_{+}^{(i)}\rangle/\sqrt{J(J+1)}|$ and (b) the momentum operator $\langle\hat{\pi}\rangle=\beta\langle\hat{J_{z}}^{(i)}\rangle/\kappa a$ as a function of the site $i$ along the chain. (c) Vertex-vertex connected correlation function $\langle\hat{J}_{+}^{(i)}\hat{J}_{-}^{(j)}\rangle_{\text{c}}/(J(J+1))=[\langle\hat{J}_{+}^{(i)}\hat{J}_{-}^{(j)}\rangle-\langle\hat{J}_{+}^{(i)}\rangle\langle\hat{J}_{-}^{(j)}\rangle]/(J(J+1))$ and (d) $zz$-connected correlation function $\langle\hat{J}_{z}^{(i)}\hat{J}_{z}^{(j)}\rangle_{\text{c}}=\langle\hat{J}_{z}^{(i)}\hat{J}_{z}^{(j)}\rangle-\langle\hat{J}_{z}^{(i)}\rangle\langle\hat{J}_{z}^{(j)}\rangle$ as a function of the site $j$ for $i=51$ (half-chain). In the plot we consider $N=101,J=20,\chi=\kappa=1,V_{\text{nn}}^{\prime}=400/\pi^{2},\lambda^{\prime}_{\kappa}=6.25/\pi^{2}$.
F.2 Time evolution of the static soliton
@@
F.3 Time evolution of the moving soliton
@@
F.4 Semi-classical phenomenological model
@@ F.4.1 Theoretical model
Figure 15: Approximation with Gaussian functions. (a) Topological charge density $\rho(x)$ of the classical soliton at $t=0$ Eq. (48) (solid blue line) and comparison with the Gaussian approximation Eq. (50) (dashed orange line). (b) $1-\cos(\beta\hat{\varphi})$ of the classical soliton at $t=0$ Eq. (49) (solid blue line) and comparison with the Gaussian approximation Eq. (51) (dashed orange line). In both plots we rescale the space by the mass term $M_{0}$. (c) Cosine interaction term $1-\cos(\beta\hat{\varphi})$ of the semi-classical static soliton ($v=0$) at $N=101,J=20,\beta^{2}=\pi/20,M_{0}^{\prime}=0.2,t=44.44$: gray dots represent the numerical results, the dash-dotted blue line the fit of the data with the Gaussian function Eq. (51). (Inset) Gaussian width $\sigma_{\text{cos}}(t)$: the solid blue line represents the width extrapolated from the previous Gaussian fit, the orange dashed line is an additional fit of the obtained width with the function $\sigma(t)=\sqrt{\sigma_{0}^{2}+(v_{\text{spr}}\cdot t)^{2}}$. (d) For the same parameters, cosine interaction term $1-\cos(\beta\hat{\varphi})$ for the semi-classical moving soliton ($v=0.3$): gray dots represent the numerical results, the dash-dotted blue line the fit with the Gaussian function Eq. (51), the dashed orange resp. the solid black line the Gaussian function in Eq. (59) neglecting resp. considering the position-dependent shift $\bar{x}\sigma_{v}v\gamma$ of the width Eq. (60). For the last two lines we use the Gaussian fit parameters extracted from the static soliton in (c).
@@
Upon neglecting the additional contribution $-\bar{x}\sigma_{v}v\gamma$, which is relevant only at short time and large distances $\bar{x}$, we retrieve
@@
F.4.2 Numerical results
We now compare the expected theoretical dependence of the observable $1-\cos(\beta\hat{\varphi})$ with the numerical results for both the semi-classical static and the moving soliton (we expect similar results for the topological charge density $\rho$). We first consider the static soliton by fitting in Fig. 15(c) the numerical data with a Gaussian distribution $p_{1}e^{-(x-p_{3})^{2}/(2p_{2})^{2}}/(\sqrt{2\pi}p_{2})$ resembling Eq. (57). We observe a good agreement of the time-dependent width $p_{2}(t)$ [see inset of Fig. 15(c)] with the theoretical predicted form $p_{2}(t)=\sqrt{(p_{2}^{1})^{2}+(p_{2}^{2}\cdot t)^{2}}$, from which the parameters $p_{2}^{1},p_{2}^{2}$ can be numerically extrapolated.
@@ F.4.2 Numerical results
F.5 Velocity fluctuations from ground-state quantum fluctuations
@@
We further recall, as discussed in App. F.1, that the two-point connected correlators of the static and moving soliton are completely determined by the ground state of the quantum sine-Gordon model.
@@
$\frac{2\pi}{\beta}\sigma(v)$ $=\sigma\left(\int_{-\infty}^{\infty}{\rm d}x\;\hat{\pi}\right)$ (65)
@@
According to the semi-classical phenomenological model [see App. F.4], the spreading velocity equals the standard deviation of the velocity $v_{\text{spr}}=\sigma(v)$, and thus we get a qualitative agreement with the spreading velocity $v_{\text{spr}}\approx 0.273\beta$ determined in the main text (the prefactors differ by a factor $\approx 1.3$). We expect the same argument to hold at $\kappa>1$, with the sum in Eq. (70) getting an additional $\kappa^{2}$ factor, which cancels out the $\kappa^{2}$ factor in Eq. (68).
@@
Appendix G Soliton-antisoliton scattering. Error estimation of the position shift
@@
More precisely, for each of the 4 regions we fix a maximal time interval where the quasi-particles move linearly, fit increasingly smaller subsets of it, sort the fit results according to the value of the slope, and choose as best guess for the fit parameters the pair containing the median value of the slope distribution. To determine the uncertainty of each fit parameter (slope and offset), we instead sort all the obtained results increasingly and estimate the error as half of the difference between the first value located at more than 0.8415 of the whole distribution and the first value at more than 0.1585 [see App. C].

### v2

@@ F.2 $\beta^{2}$-renormalization and $\kappa$-dependence
To engineer the spatial derivative term $\partial_{x}\hat{\varphi}(x)]^{2}$ in the sG model, we need to Taylor expand the cosine term $V_{\text{nn}}^{\prime}/2\cdot\cos(\varphi_{i}-\varphi_{i+1})=-V_{\text{nn}}^{\prime}/2\cdot\cos(\delta\varphi_{i})$ of the lattice QFT Hamiltonian Eq. (38), which arises from the dipole-dipole interaction term $\frac{1}{4}\left(\hat{J}_{+}^{(i)}\hat{J}_{-}^{(j)}+\mathrm{H.c.}\right)$. We thus obtain
@@
where we need to neglect higher-order term $(n\geq 2)$ to retrieve the sG model. Upon multiplying the Hamiltonian with $\beta^{2}/(2\chi\kappa^{2}a)$, mapping $\kappa\hat{\varphi}_{\text{latt}}\rightarrow\beta\hat{\varphi}_{\text{cont}}$, and using the relation $V_{\text{nn}}^{\prime}/\chi=4\kappa^{4}/\beta^{4}$, as we do in App. C to obtain the continuum sG model, the neglected terms become
@@ F.3 $M_{0}^{\prime}$ renormalization
F.4 Lukyanov-Zamolodchikov conjecture
@@
with $K=\beta^{2}/8\pi$ and is valid for $\beta^{2}<8\pi$, $|\Re\gamma|<\sqrt{8\pi}/(2\beta)$. For $\gamma=\sqrt{K}=\beta/\sqrt{8\pi}$ the above expression agrees with the theoretical prediction for the vertex operator of the sG model Eq. (14) [48].
Appendix G Semi-classical soliton preparation and dynamics
In this Appendix we provide additional details regarding the preparation of the semi-classical soliton, as well as its dynamics. This Appendix is structured as follows: in App. G.1 we discuss the prepared soliton state, and we show that the two-point correlators remain untouched, in App. G.2 resp. App. G.3 we focus on the time evolution of the static resp. moving soliton, in App. G.4 we discuss in more detail the semi-classical model, while in App. G.5 we illustrate the relation between the $zz$-correlator of the quantum ground state and the standard deviation of the velocity of the semi-classical model.
G.1 State preparation and correlation functions
@@
[…]
@@
Figure 12: One- and two-point correlation functions of the prepared soliton state. Comparison of selected expectation values for different initial states: the solid black line represents the ground state of the quantum sG model, the blue circles the semi-classical static soliton $v=0$, the orange dots the semi-classical moving soliton at $v=0.5$, the dashed black line the theoretical prediction for the momentum of the classical moving soliton. (a) Absolute value of the expectation value of the vertex operator $|\langle\hat{J}_{+}^{(i)}\rangle/\sqrt{J(J+1)}|$ and (b) the momentum operator $\langle\hat{\pi}\rangle=\beta\langle\hat{J_{z}}^{(i)}\rangle/\kappa a$ as a function of the site $i$ along the chain. (c) Vertex-vertex connected correlation function $\langle\hat{J}_{+}^{(i)}\hat{J}_{-}^{(j)}\rangle_{\text{c}}/(J(J+1))=[\langle\hat{J}_{+}^{(i)}\hat{J}_{-}^{(j)}\rangle-\langle\hat{J}_{+}^{(i)}\rangle\langle\hat{J}_{-}^{(j)}\rangle]/(J(J+1))$ and (d) zzzz-connected correlation function $\langle\hat{J}_{z}^{(i)}\hat{J}_{z}^{(j)}\rangle_{\text{c}}=\langle\hat{J}_{z}^{(i)}\hat{J}_{z}^{(j)}\rangle-\langle\hat{J}_{z}^{(i)}\rangle\langle\hat{J}_{z}^{(j)}\rangle$ as a function of the site $j$ for $i=51$ (half-chain). In the plot we consider $N=101,J=20,\chi=\kappa=1,V_{\text{nn}}^{\prime}=400/\pi^{2},\lambda^{\prime}_{\kappa}=6.25/\pi^{2}$.
G.2 Time evolution of the static soliton
@@
G.3 Time evolution of the moving soliton
@@
G.4 Semi-classical phenomenological model
@@ G.4.1 Theoretical model
[…]
@@
Upon neglecting the additional contribution $\bar{x}\sigma_{v}v\gamma$, which is relevant only at short time and large distances $\bar{x}$, we retrieve
@@
G.4.2 Numerical results
[…]
@@ G.4.2 Numerical results
G.5 Velocity fluctuations from ground-state quantum fluctuations
@@
We further recall, as discussed in App. G.1, that the two-point connected correlators of the static and moving soliton are completely determined by the ground state of the quantum sine-Gordon model.
@@
[…]
@@
According to the semi-classical phenomenological model [see App. G.4], the spreading velocity equals the standard deviation of the velocity $v_{\text{spr}}=\sigma(v)$, and thus we get a qualitative agreement with the spreading velocity $v_{\text{spr}}\approx 0.273\beta$ determined in the main text (the prefactors differ by a factor $\approx 1.3$). We expect the same argument to hold at $\kappa>1$, with the sum in Eq. (70) getting an additional $\kappa^{2}$ factor, which cancels out the $\kappa^{2}$ factor in Eq. (68).
@@
Appendix H Soliton-antisoliton scattering. Error estimation of the position shift
@@
More precisely, for each of the 4 regions we fix a maximal time interval where the quasi-particles move linearly, fit increasingly smaller subsets of it, sort the fit results according to the value of the slope, and choose as best guess for the fit parameters the pair containing the median value of the slope distribution. To determine the uncertainty of each fit parameter (slope and offset), we instead sort all the obtained results increasingly and estimate the error as half of the difference between the first value located at more than 0.8415 of the whole distribution and the first value at more than 0.1585 [see App. D].